Each side of a square is increased 7 inches. When this happens, the area is multiplied by 4. How many inches in the side of the original square?
Answer #1

\blacksquare\,\,\,Let\,\,the\,\,original\,\,length\,\,of\,\, each\,\,side\,\,of\,\,the\,\,square\\is\,\,x\,\,inches\,\,.\\\\Now,\,\, according\,\,to\,\,the\,\, question\,\,if\,\,each\,\,side\,\,of\,\,the\,\,square\\is\,\, increased\,\,7\,\,inches,\,\,the\,\,area\,\, is\,\, multiplied\,\,by\,\,4\,\,\\\\\therefore\,\,4x^{2}=(x+7)^{2}\\\\\Rightarrow\,\, 4x^{2}=x^{2}+2\times x\times 7+7^{2}\\\\\Rightarrow\,4x^{2}=x^{2}+14x+49\\\\ \Rightarrow\,\,4x^{2} -x^{2}-14x-49=0\\\\\Rightarrow\,\,3x^{2}-14x-49=0 \\\\\Rightarrow\,\,3x^{2}-21x+7x-49=0\\\\\Rightarrow\,3x(x-7)+7(x-7)=0\\\\\Rightarrow \,(x-7)(3x+7)=0\\\\Either,\,\,(x-7)=0\,\,\,\,\,\,\,\,\,|\,\,\,\,\,or,(3x+7)=0\\\\ \Rightarrow \,x=7\hspace{74pt}|\,\,\Rightarrow\,x=\frac {-7}{3}\\\\Since, \,\,the\,\,length\,\,can\,\, not\,\, negative\,\,so,\,\,length\,\,each\,\,side\\of\,\, the\,\,square\,\,is,\,\,7\,\,inches\,.

••••••••••••••••

[ Hope you understand.....for any further assistance or clarification please comment.....!!!]